OpenStudy (anonymous):

Medal for who ever helps me! A coffee shop pays Coffee Delivery Company A, a certain price for each disposable cup it orders plus a weekly delivery charge to remain on the driver’s delivery route. The cups are purchased in increments of 500. To quickly determine how much the coffee shop will be spending on cups before their arrival, the owner created the following table:

OpenStudy (anonymous):

Cups Ordered Price 0 $7 500 $25.50 1,000 $44 1,500 $62.50 2,000 $81 2,500 $99.50 3,000 $118

OpenStudy (anonymous):

A.The price of the cups, p(x), is a function of the number of x cups ordered. Using the table, determine the average rate of change for the first 1,000 cups ordered and then for all 3,000 cups ordered. What does this tell you about the function? B. Use the table to evaluate and interpret p(0). What is a possible explanation for this? C. Sketch a graph, labeling its key features, to show the price the coffee shop would pay Coffee Delivery A to have between 0 and 3,000 cups delivered each week. D.Create a model using function notation that represents how the two quantities, cups and cost, are related. E. The coffee shop found another delivery company that sells orders at increments of 500 cups, Coffee Delivery B. They charge $3.50 each week to be on their delivery route and charge 3.9 cents per disposable cup.

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@kirbykirby

OpenStudy (imstuck):

Sorry. I was out. Let me look at it; it appears that you still need help?

OpenStudy (imstuck):

For a, the average rates of change are the same. In order to find the average rate of change you subtract y2 - y1/x2 - x1. y2 is 44, y1 is 7, x2 is 10000 and x1 is 0. That rate of change gives you 44-7/1000-0=37/1000 = .037. For all 3000 cups, you have the same operation. 118-7/3000-0=111/3000=.037. That means that for every increment of 500 cups you are paying the same for every 500 cups. It doesn't save you money to buy in bulk, obviously!

OpenStudy (anonymous):

Thanks man. That's really all I need help with.